OpenStudy (anonymous):

8* 1/8 + (-1) + 5 =5 Select all that justify the following statement. commutative - addition inverse - addition associative - multiplication symmetric commutative - multiplication associative - addition inverse - multiplication

OpenStudy (anonymous):

\[8\times \frac{1}{8}=1\] inverse multiplication \[1 (-1)=0\] inverse addtion

OpenStudy (anonymous):

sorry i meant \[1+(-1)=0\]